Telescope Bootstrap Startup: 13 September 2012 UT

Observer: JHill (LBTO), OKuhn(Tucson)
Telescope Operator: GBechetti (LBTO)
Software Support: CBiddick (Tucson)
AO Support: DMiller (Tucson)
Telescope Support: MMidkiff (LBTO)
Instrument Support: ESolheid (LBTO)

Summary

Sky was clear. The first half had heavy fog. The second half had less fog, but still large quantities of water dripping off the dome.

We did lots of closed dome testing: more MODS1 calibrrations by Olga, LBTI DX AOS testing by Vidhya, and general TCS testing of 2012G by John and Geno.

Issue 4214: GSC can't communicate with AGW1 and AGW2 for homing.

Preparations

MODS1 AGw Check

I've used GCS to read the guider and wfs images from MODS. They look fine in the dark with the hatch closed. The images are stored in /home/LBTO/AGW_Data/20120912_MODS/. (The traditional location of /OldRepository/AGW_Data is not writeable.)

Weather

We've seen the sun, and our vitamin D levels are stimulated. Thus, we are removing the retroreflector and configuring the telescope for MODS starting at 1:00 UT. Note however, that we are still very close to the dewpoint, so it could be a long foggy night. At 00:00 UT, the are thunderstorms in San Carlos and in Redington moving ESE. It looks like one with go south of us, and the other north.

Doug fixed IDL in the LBTO account, and added LBTIR and LBTIL as instrument options.

Details

Things are never dull here

01:30 The crane scale battery dies while removing the retroreflector.

01:35 While looking for the battery, Elliot discovers a pool of glycol (~5 gallons) that has leaked out of the LUCI chiller next to the bell jar. The leak is cause by a cracked fitting. The rate of the leak is about that of a hypoderimic needle, so it has probably been leaking all day. Leak management and retroremoval continue in parallel.

02:25 The fog bank has caught up to us. Humidity is 98%, although the sky above is clear.

03:55 All the convective activity in southern Arizona has ended, but here we are sitting in a fog bank with 99% humidty. T=7 degC.

More MODS Calibrations

04:00 At zenith with dome dark for Olga to take more MODS1 calibrations.
  • Finishing up the set of calibrations not done last night. mods1?.20120913.nnnn.fits
    • doBias (2-11),
    • prbias (12-16),
    • prflats_b.cal (b 17-21 Vflat 5 ~40k counts max w clear, 22-26 ~35 counts max w/ UG5),
    • prflats_r.cal (r 17-21 Vflat 4 GG495 ~30k counts max),
    • prlamps_b.cal (b 27 ok,28 ok,)29 Kr and 30 Xe both have very saturated lines at red end),
    • prlamps_r.cal (r 22-25)
    • prlamps.cal (b 31-34 , r 26-29) and
    • imflats.cal
      • timeout on calmode, retry OK
      • b 35-44
        • u_sdss using QTH2 is saturated!
        • g_sdss w VFLAT 4 44k,50k,..57k --- looks like VFLAT gets brighter with time since turning it on (see mods1b.20120913.0040-0044.fits) VFLAT was off for ~50 minutes prior to this;
      • r 30- 42 imflats.cal i_sdss and z_sdss flats saturated.
        • r_sdss w VFLAT 2 has 60k counts max, i_sdss is completely saturated with VFLAT 1.5 and 15s, z_sdss at VFLAT 1.5 and 8s is also saturated - aborted sequence before script ended), imflats_b.cal, imflats_r look like they use the same settings as imflats.cal so I will not do these.
        • manually take z_sdss flat with 2 seconds VFLAT 2 ---> #43 (60k counts max)
        • VFLAT 1 and 2 seconds ---> #44 (3000 counts max).
        • VFLAT 2 and 2 seconds ---> #45 (65k max)
        • VFLAT 2 and 1.5 seconds ---> #46 (52k max)
        • manually take i_sdss flat with VFLAT 2 and 1.5 seconds ---> #47 (30k) 4 more: 48-51.
        • four more z_sdss flats with VFLAT 2 and 2 seconds ---> 52-54...
        • four more z_sdss flats with VFLAT 2 and 1.5 seconds ---> 55-58.
  • Collimator TTF need to be entered still, but they did not change much.
  • Plan tonight for MODS - if we open - is to observe a Stone field to check alignment (StoneO_1136 has a good RA for the start of the night). Scripts are in /home/LBTO/kuhn/MODS/StartupScripts/AlignCheck/
    • First just one dual imaging snapshot of 30-seconds, StoneO_1136_act0.img
    • Next four track presets to the same RA,DEC but different PAs: StoneO_1136_trk0.img, StoneO_1136_trk180.img and StoneO_1136_trk90.img and StoneO_1136_trk270.img (in pairs separated by 180 degrees since the IRAF rotcenter task requires this).
    • Finally, two scripts do 5-dice: StoneO_1136_act0_5dicexy.img uses offsetxy x y abs to make absolute detector offsets and StoneO_1136_act0_5dice.img uses offset x y abs to make absolute RA,DEC offsets. The grid spacing is 10 arcseconds for each.

04:10 Open vent doors partway to cool the chamber. Still foggy, but the chamber is 8 degC above the dewpoint.

06:30 After brief drop in humidity we are back up to 98%.

Authorize LUCI on both sides Turn on LFBG/RFBG rotators Turn on AGW1 and AGW2 - AGW2 gets an RPC error.

LBTI AO tests on DX

06:45 Vidhya is sending test LBTI presets while Geno exercises the AGWs.

07:00 reboot agw-control to restart oacserver

AGW selection error for LUCIs

08:03 failure of AGW selection when authorizing LUCI on both sides.

Stop and restart both GCSs

08:08 Same failure again

08:09 Authorize MODS/None and AGW selection succeeds.

08:11 Authorize LUCI/None makes the same failure.

08:15 Power off both AGWs. Stop GCSs. Reboot agw-control.

Authorize LUCI/LUCI repeats the same failure. tcslog says the error is at the homing of the stages. Geno gets "Network Error" when he tries to home probe by hand.

08:30 Power off both AGWs, reboot oacontrol, but do NOT restart GCS

08:30 Resuming MODS calibrations: Edited version of imflats_r.cal to reduce exptime for r to 15 sec (25k counts), and use VFLAT 2 and exptime 1.5 for i and z. Note these are red-only so not quite the same as the ones on which the scaling was made earlier.
  • VFLAT brightening during first (r_sdss mods1r #59-63) sequence

08:38 now the manual home command works

08:41 start GCSs. Now AGW works OK on the left side.

AGW Pupil Check

AGW1 @ LFBG

Images are in /home/lbto/AGW_Data/20120913_AGW1

Low Lights on

wfs 0001 - SX M3 was not in place

AGW2 @ RFBG

Images are in /home/lbto/AGW_Data/20120913_AGW2

08:53 DX M3 is at the Andrew/Lee/Bob position of 20120904.

wfs 0001 - 0 deg wfs 0002 - 90 deg wfs 0003 - 180 deg wfs 0004 - 270 deg

Oops, the above were with the probe at the home/park position at the edge of the field.

AGW1 @ LFBG

09:02 back to SX: setGSsfp left to put probe in center.

wfs 0005 - 0 deg 10 sec wfs 0007 - 90 deg wfs 0008 - 180 deg wfs 0009 - 270 deg

AGW2 @ LFBG

09:22 setGSsfp right

wfs 0005 - 0 deg 10 sec wfs 0006 - 90 deg wfs 0007 - 180 deg wfs 0008 - 270 deg

Rotator Speeds

Geno measures time for a 360 rotation of the rotator

RFBG: 105 sec LFBG: 6 min LDG: 105 sec

Change from 2012G to 2012F

09:40 start switching

09:52 running 2012F (had a problem with RPC Server on obs2)

09:54 authorizing LUCI/None succedes.

09:55 authorizing LUCI/LUCI succedes.

some other error on LBTI/LBTI

10:07 authorizing LUCI/LUCI succedes.

10:09 authorizing LBTI/LBTI succedes

11:00 Hygrometer on the roof reads 74%, but water is dripping everywhere off the dome.

-- JohnHill - 13 Sep 2012
Topic revision: r9 - 15 Sep 2012, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback